OpenStudy (luigi0210):

For y = -x^4+2x^2+2 on the interval [-3, 2] what would be the absolute min/max and local min/max?

OpenStudy (anonymous):

differentiate the function . set the derivative equal to zero solve for x to get critical points .

OpenStudy (luigi0210):

Well I know the derivative would be 4x^3+4x And the critical points are -1, 0, and 1

OpenStudy (luigi0210):

 and this would be the sign graph right?

OpenStudy (anonymous):

you sure these are the critical points ?

OpenStudy (luigi0210):

yea -4x^3+4x -4x(x^2-1)=0 ------------------ -4x=0 x^2-1=0 ------------------ x=0/-4=0 (x+1)(x-1)=0 ------------------ x=0, -1, 1

OpenStudy (anonymous):

ok , i was confused when you said above derivative is 4x^3+4x !!! ok f'(x)=-4x^3+4x -4x^3+4x=0 x=0,-1,1 evaulate the function f(x) at the critical points .let me know what you get

OpenStudy (anonymous):

for example evaulating the function at critical point 0 put x=0 in the given function y=f(0)=(0)^4+2(0)^2+2=2 similarly find f(1) and f(-1).

OpenStudy (luigi0210):

I don't really have trouble with that kind of stuff, I just need help on telling the difference between a local ma/min and an absolute max/min

OpenStudy (luigi0210):

Like, how can you tell if a point is an abs max and a relative max or just an abs max? How can some points be both?

OpenStudy (anonymous):

ok the largest max value a function will have is its absolute maxima the lowest minmium value a function willl have its abosolute minima in the given function evaluate the function at x=0 ,1,-1 also at the end points x=-3 and x=2 so f(0)=2 f(1)=5 f(-1)=5 f(-3)=105 f(2)=26 the largest value is at x=-3 105 which is absolute maxima of the function the smallest value is at x=0 which is the absolute minima of the function .

OpenStudy (luigi0210):

Oh, well I could agree with on the max but I had the same thing you have for min, but I got it marked wrong

OpenStudy (anonymous):

i don't know why your teacher marked it as wrong because its the smallest pssible value in the interval [-3,2] the function will take . that is 2.

OpenStudy (luigi0210):

I don't know either.. but she hasn't been wrong yet

OpenStudy (anonymous):

@satellite73

OpenStudy (luigi0210):

Would you use the second derivative test for anything on this?

OpenStudy (anonymous):

check them

OpenStudy (anonymous):

you are looking for the absolute max and min check the critical points, i.e. plug them in check the endpoints of the interval the biggest is the max and the smallest is the min

OpenStudy (anonymous):

however i think \(f(1)=3\)

OpenStudy (anonymous):

also \(f(-1)=3\)

OpenStudy (anonymous):

it is \( f(x) = -x^4+2x^2+2\) right?

OpenStudy (anonymous):

mistake i think was using \(x^4+2x^2+2\) and dropping the minus sign

OpenStudy (anonymous):

yup i missed minus sign.